# In Search of Lost Online Test-time Adaptation: A Survey

### v2

@@ 3.1.1 Normalization Calibration
Motivation. In domain adaptation, aligning batch normalization statistics is shown to mitigate performance degradation brought by covariate shifts. The underlying hypothesis suggests that information about labels is encoded within the weight matrices of each layer. At the same time, knowledge related to specific domains is conveyed through the statistics of BatchNorm layer. Consequently, an adaptation based on updating BatchNorm layer could bring enhanced performance of unseen domains (Li et al, 2017). A similar idea can be used in online test-time adaptation. The assumption is, given a neural network $f$ trained on a source dataset $\mathcal{D}_{S}$ with normalization parameters $\beta$ and $\gamma$, updating $\{\gamma,\beta\}$ based on test data $\bm{x}_{i}^{t}$ at each time step $t$ will improve $f$’s robustness on the test domain. Building upon the assumption, initial investigations in OTTA predominantly revolved around fine-tuning only on updating the normalization layers. This strategy has a few popular variations. A common practice adjusts statistics ($\mu$ and $\sigma$) and affine parameters ($\beta$ and $\gamma$) in the BatchNorm layer. Note that the choice of normalization techniques, such as LayerNorm or GroupNorm, may depend on the backbone architecture and specific optimization objectives.
Tent (Wang et al, 2021a) and its subsequent works, such as (Niu et al, 2022; Jang et al, 2023), are representative approaches within this paradigm. They update the statistics and affine parameters of BatchNorm for each test batch while freezing the remaining parameters. However, as seen in Tent, updated by minimizing soft entropy, the effectiveness of batch-level updates is dependent on data quality within each batch, introducing potential performance fluctuations. For example, noisy or poisoned data with biased statistics would significantly influence the BatchNorm updates. Methods that aim at stabilization via dataset-level estimate are proposed to mitigate such performance fluctuation arising from batch-level statistics. Gradient preserving batch normalization GpreBN (Yang et al, 2022) is introduced that allows for cross instance gradient backpropagation by modifying the BatchNorm normalization factor:
@@ 3.1.3 Optimization Objective
Optimizing (increasing) confidence. Covariate shifts typically lead to lower model accuracy, which in turn causes the model to express high uncertainty. The latter is often observed As such, to improve model performance, an intuitive way is to enhance model confidence for the test data.
@@
Tent is a popular method that uses entropy minimization to update the affine parameters in BatchNorm. Subsequent studies have adopted and expanded upon this strategy. For example, EATA (Niu et al, 2022) implements a sample selection strategy on top of entropy minimization by only updating reliable and unique samples. Following EATA, Seto et al (2023) introduces entropy minimization by incorporating self-paced learning. This addition ensures the learning process progresses at an adaptive and optimal pace. Furthermore, by integrating the general adaptive robust loss (Barron, 2019) into self-paced learning, the proposed method achieves robustness against large and unstable loss values. TTPR (Sivaprasad et al, 2021) combines entropy minimization with prediction reliability, which operates across various views of a test image to form a consistency loss. This is done by merging the mean prediction across three augmented versions, with each augmented prediction separately. Lin et al (2023) introduces minimizing the entropy loss on augmentation-averaged predictions while assigning high weights for low-entropy samples. In SAR (Niu et al, 2023), when minimizing entropy, an optimization strategy is used so that parameters enabling ‘flatter’ minimum regions can be found. This is shown to allow for a better model updating stability.
@@
A drawback of entropy minimization is the overly early convergence because gradients of high-confidence predictions do not contribute much. To allow contributions from reliable predictions and thus smooth updating, SLR (Mummadi et al, 2021) borrows the negative log-likelihood ratio NLLR(Zhu et al, 2018)
$R\left(\hat{y},y^{r}\right)=-\sum_{c}y_{c}^{r}\log\frac{\hat{y}_{c}}{\sum_{i\neq c}\hat{y}_{i}}$ (6)
$=H\left(\hat{y},y^{r}\right)+\sum_{c}y_{c}^{r}\log\sum_{i\neq c}\hat{y}_{i},$
to design their loss. Here, $\hat{y}$ is the prediction that wants to approach $y^{r}$, and $c$ is the predicted class. The entropy $H$ is lower bound by 00. Even if $y_{c}^{r}\rightarrow 1$ and $\sum_{i\neq c}\hat{y}_{i}\rightarrow 0$, where traditional entropy minimization occurs the gradient vanishing issue for high-confidence predictions, NLLR could induce non-vanishing gradients since $\log\sum_{i\neq c}\hat{y}_{i}\rightarrow-\infty$.
Another drawback of entropy minimization is that it might obtain a degenerate solution where every data point is assigned to the same class. To avoid this, MuSLA (Kingetsu et al, 2022) employs mutual information of the sample $X$ and the corresponding prediction $\hat{Y}$,
@@
The cooperation between a teacher and a student is another possible solution to optimize prediction confidence with reliability. Here, the teacher is usually the moving average of the model of interest across iterations. CoTTA (Wang et al, 2022a) uses the Softmax prediction from the teacher to supervise the Softmax predictions from the student under the cross-entropy loss.
@@
A cross-entropy loss helps minimize model prediction uncertainty but may fail to give consistent uncertainty scores under different augmentations, model randomness (MC dropout), etc. This problem is more critical when the teacher-student mechanism is not used. To address this problem, MEMO (Zhang et al, 2022) computes the average prediction across multiple augmentations for each test sample and then minimizes the entropy of the marginal output distribution over augmentations:
@@
This term aims to identify a weight perturbation $\epsilon$ within a Euclidean ball of radius $\rho$ that maximizes entropy. It quantifies sharpness by measuring the maximal change in entropy between $\Theta$ and $\Theta+\rho$. As such, Eq. 11 jointly minimizing entropy and its sharpness. Gong et al (2023) uses same idea for their optimization.
@@ 3.2.1 Data Augmentation
To ensure consistent and reliable predictions, PAD (Wu et al, 2021) employs multiple augmentations of a single test sample for majority voting. This is grounded in the belief that if the majority of the augmented views yield the same prediction, it is likely to be correct, as it demonstrates insensitivity to variations in style. Instead, TTPR (Sivaprasad et al, 2021) adopts KL divergence to achieve consistent predictions. For every test sample, it generates three augmented versions. The model is then refined by aligning the average prediction across these augmented views with the prediction for each view. Another approach is MEMO, which uses AugMix (Hendrycks et al, 2020) for test images. For a test data point, a range (usually $32$ or $64$) of augmentations from the AugMix pool $\mathcal{A}$ is generated to make consistent predictions.
@@ 3.2.2 Memory Bank
Contrastive learning is well-suited for OTTA, as discussed in 3.1.3. However, this approach can be challenging for online learning, especially when it pushes away feature representations of data from the same class. Unlike conventional methods where the feature space can be revisited multiple times, AdaContrast (Chen et al, 2022) offers an innovative solution. It keeps all previously encountered key features and pseudo-labels in a memory queue to avoid forming ’push-away’ pairs from the same class. This method speeds up the learning process and reduces the risk of error accumulation in data from the same class, thereby improving the efficiency and precision of the learning process.
@@ 3.3.1 Module Addition
In the case of continual adaptation, promptly detecting and adapting to changes in data distribution is inevitable to deal with catastrophic forgetting and the accumulation of errors. To realize this, ViDA (Liu et al, 2023b) utilizes the idea of low/high-rank feature cooperation. Low-rank features retain general knowledge, while high-rank features better capture distribution changes. To obtain these features, the authors introduce two adapter modules correspondingly parallel to the linear layers (if the backbone model is ViT). Additionally, since distribution changes in continual OTTA are unpredictable, strategically combining low/high-rank information is crucial. Here, the authors use MC dropout (Gal and Ghahramani, 2016) to assess model prediction uncertainty about input $x$. This uncertainty is then used to adjust the weight given to each feature. Intuitively, if the model is uncertain about a sample, the weight of domain-specific knowledge (high-rank feature) is increased, and conversely, the weight of domain-shared knowledge (low-rank feature) is increased. This helps the model dynamically recognize distribution changes while preserving its decision-making capabilities.
@@ 3.3.2 Module Substitution
Layer substitution typically refers to swapping an existing layer in a model with a new one. The commonly used techniques are about:
@@ 3.3.3 Prompt-based Method
The rise of vision-language models demonstrates their remarkable ability in zero-shot generalization. However, these models often underperform for domain-specific data. While attempting to address this, traditional fine-tuning strategies typically compromise the model’s generalization power by altering its parameters.
In response, borrows the idea from test-time adaptation, Test Time Prompt Tuning emerges as a solution. Deviating from the conventional methods, it fine-tunes the prompt, adjusting only the context of the model’s input, thus preserving the generalization power of the model. One representative is TPT (Shu et al, 2022). It generates $N$ randomly augmented views of each test image and updates the prompting parameter by minimizing the entropy of the averaged prediction probability distribution. Additionally, a confidence selection strategy is proposed to filter out the output with high entropy to avoid noisy updating brought by unconfident samples. By updating the learnable parameter of the prompt, it could be easier to adapt the model to the new, unseen domains.
(a) CIFAR-10-C &\&& 10.1 Dataset
(b) ImageNet-C Dataset
(c) CIFAR-10-Warehouse Dataset
@@
The prompt-related ideas are also powerful in OTTA tasks. “Decorate the Newcomers” (DN) (Gan et al, 2023) employs prompts as supplementary information added onto the image input. To infuse the prompts with relevant information, it employs a student-teacher framework in conjunction with a frozen source pre-trained model to capture both domain-specific and domain-agnostic prompts. For acquiring domain-specific knowledge, it optimizes the cross-entropy loss between the outputs of the teacher and student models. Additionally, DN introduces a parameter insensitivity loss to mitigate the impact of parameters prone to domain shifts. This strategy aims to ensure that the updated parameters, which are less sensitive to domain variations, effectively retain domain-agnostic knowledge. Through this approach, DN balances learning new, domain-specific information while preserving crucial, domain-general knowledge.
@@ 3.3.4 Summary
4 Empirical Study
In this empirical study, we focus on upgrading existing OTTA methods for the Vision Transformer (ViT) model (Dosovitskiy et al, 2021), investigating their potential to migrate to new generations of backbones. We provide solutions for adapting methods originally proposed for the CNN architecture to ViTs.
Baselines. We benchmark seven OTTA methods. To ensure fairness, we adhere to a standardized testing protocol and select five datasets, including three corrupted ones (i.e., CIFAR-10-C, CIFAR-100-C, and ImageNet-C), one real-world shifted dataset (CIFAR-10.1), and one comprehensive dataset (CIFAR-10-Warehouse). The CIFAR-10-Warehouse repository plays a crucial role in our evaluation, offering a wide range of subsets, including real-world variations sourced from different search engines and images generated through diffusion processes. Specifically, our survey focuses on two subsets of the CIFAR-10-Warehouse dataset: the Google split and the Diffusion split. These subsets, embodying both real-world and artificial data shifts, facilitate a comprehensive assessment of OTTA methods.
@@ 4.1 Implementation Details
Optimization details. We employ PyTorch for implementation. The foundational backbone for all approaches is ViT-base-patch16-224 (Dosovitskiy et al, 2021) [^1] https://github.com/huggingface/pytorch-image-models. When using CIFAR-10-C, CIFAR-10.1, and CIFAR-10-Warehouse as target domains, we train the source model on CIFAR-10 with $8,000$ iterations, including a warm-up phase spanning 1,600 iterations. The training uses a batch size 64 and the stochastic gradient descent (SGD) algorithm with a learning rate of $23e-$ 2. We mostly use an identical configuration to train the source model on CIFAR-100, with an extended training duration of $16,000$ iterations and a warm-up period spanning $4,000$ iterations. The source model on the ImageNet-1k dataset is acquired from the Timm repository 222vit_base_patch16_224.orig_in21k_ft_in1k. Additionally, we apply basic data augmentation techniques, including random resizing and cropping, across all methods. The Adam optimizer with a momentum term $\beta$ of $0.9$ and a learning rate of $31e-$ 3 ensures consistency during adaptation. Resizing and cropping techniques are applied as a default preprocessing step for all datasets. Then, a uniform normalization $(0.5,0.5,0.5)$ is adopted to mitigate potential performance fluctuations arising from external factors beyond the algorithm’s core operations.
@@
Figure 8: Comparison of the OTTA performance on the CIFAR-10-C (at severity level $5$) and CIFAR-10.1 datasets. The upper and bottom plots show the experiments conducted with batch size $16$ and $1$ based on the LayerNorm updating strategy, respectively.
@@
Component substitution. To successfully adapt core methods for use with the Vision Transformer (ViT), we have developed a series of strategies:
  • Switch to LayerNorm: In light of the absence of BatchNorm layer in ViT, we substitute all BatchNorm updates with LayerNorm updates.
@@
  • Disregard BatchNorm mixup: Removing statistic mixup strategy originally designed for BatchNorm-based methods as LayerNorm is designed to normalize each data point independently.
@@
  • Sample Embedding Changes: For OTTA methods that rely on feature representations, an effective solution is to use the class embedding (i.e., the first dimension of the ViT feature) as the image feature.
@@
  • Pruning Incompatible Components: Any elements incongruent with the ViT framework should be identified and removed.
@@
These strategies lay the foundation for integrating core OTTA methods with the Vision Transformer, thereby broadening their application to this advanced model architecture. It’s important to note that these solutions are not exclusively limited to the OTTA methods. Rather, they can be viewed as a broader set of guidelines that can be applied where there is a need for upgrading to a new generation of backbone architectures.
Baselines: We carefully select seven methods to examine the adaptability of OTTA methods thoroughly. They include:
@@
    Tent: A fundamental OTTA method rooted in BatchNorm updates. To reproduce it on ViTs, we replace its BatchNorm updates with a LayerNorm updating strategy.
@@
    CoTTA employs the mean-teacher model, parameter reset, and selective augmentation strategy. While it necessitates updating the entire student network, we further assess the LayerNorm updating strategy on its student model. We also deconstruct its parameter reset strategy, resulting in four variants: parameter reset with LayerNorm updating (CoTTA-LN), parameter reset with full network updating (CoTTA-ALL), updating LayerNorm without parameter reset ($*{}^{*}$-LN), and full network updating without parameter reset ($*{}^{*}$-ALL).
@@
    SAR follows the same strategy as Tent while using sharpness-aware minimization for optimization.
@@
    Conj-PL: As the source model optimized by cross-entropy loss, it is similar to Tent but allows the model to interact with the data twice: once for updating LayerNorm and another for prediction.
@@
    MEMO: Two versions of MEMO are considered: full model updating and LayerNorm updating. We remove all data normalizations from its augmentation set to maintain consistency and prevent unexpected performance variations.
@@
    RoTTA: Due to the architecture limitation from ViTs, we exclude the RBN module in RoTTA. As LayerNorm in ViT is designed to handle data at the sample level, the RBN module is inapplicable.
@@
    TAST: We use the first dimension’s class embedding as the feature representation to suit ViT architecture.
Despite the large OTTA method pool, thoroughly evaluating this selected subset could yield valuable insights commensurate with our expectations. In the empirical study, we address the following key research questions:
4.2 Is OTTA still working with ViT?
To assess the efficacy of selected methods, we compare them against the source-only (i.e., direct inference) baseline. We discuss the experimental result for each dataset in the subsequent sections.

### v3

@@ 3.1.1 Normalization Calibration
Motivation. In domain adaptation, aligning batch normalization statistics is shown to mitigate performance degradation brought by covariate shifts. The underlying hypothesis suggests that information about labels is encoded within the weight matrices of each layer. At the same time, knowledge related to specific domains is conveyed through the statistics of BatchNorm layer. Consequently, an adaptation based on updating BatchNorm layer could bring enhanced performance of unseen domains (Li et al, 2017). A similar idea can be used in online test-time adaptation. The assumption is, given a neural network $f$ trained on a source dataset $\mathcal{D}_{S}$ with normalization parameters $\beta$ and $\gamma$, updating $\{\gamma,\beta\}$ based on test data $\bm{x}_{i}^{t}$ at each time step $t$ will improve $f$’s robustness on the test domain.
Building upon the assumption, initial investigations in OTTA predominantly revolved around fine-tuning only on updating the normalization layers. This strategy has a few popular variations. A common practice adjusts statistics ($\mu$ and $\sigma$) and affine parameters ($\beta$ and $\gamma$) in the BatchNorm layer. Note that the choice of normalization techniques, such as LayerNorm or GroupNorm, may depend on the backbone architecture and specific optimization objectives.
Tent (Wang et al, 2021a) and its subsequent works, such as (Jang et al, 2023), are representative approaches within this paradigm. They update the statistics and affine parameters of BatchNorm for each test batch while freezing the remaining parameters. However, as seen in Tent, updated by minimizing soft entropy, the effectiveness of batch-level updates is dependent on data quality within each batch, introducing potential performance fluctuations. For example, noisy or poisoned data with biased statistics would significantly influence the BatchNorm updates. Methods that aim at stabilization via dataset-level estimate are proposed to mitigate such performance fluctuation arising from batch-level statistics. Gradient preserving batch normalization GpreBN (Yang et al, 2022) is introduced that allows for cross instance gradient backpropagation by modifying the BatchNorm normalization factor:
@@ 3.1.3 Optimization Objective
Optimizing (increasing) confidence. Covariate shifts typically lead to lower model accuracy, which in turn causes the model to express high uncertainty. The latter is often observed. As such, to improve model performance, an intuitive way is to enhance model confidence for the test data.
@@
Tent is a popular method that uses entropy minimization to update the affine parameters in BatchNorm. Subsequent studies have adopted and expanded upon this strategy. For example, Seto et al (2023) introduces entropy minimization by incorporating self-paced learning. This addition ensures the learning process progresses at an adaptive and optimal pace. Furthermore, by integrating the general adaptive robust loss (Barron, 2019) into self-paced learning, the proposed method achieves robustness against large and unstable loss values. TTPR (Sivaprasad et al, 2021) combines entropy minimization with prediction reliability, which operates across various views of a test image to form a consistency loss. This is done by merging the mean prediction across three augmented versions, with each augmented prediction separately. Lin et al (2023) introduces minimizing the entropy loss on augmentation-averaged predictions while assigning high weights for low-entropy samples. In SAR (Niu et al, 2023), when minimizing entropy, an optimization strategy is used so that parameters enabling ‘flatter’ minimum regions can be found. This is shown to allow for a better model updating stability.
@@
A drawback of entropy minimization is that it might obtain a degenerate solution where every data point is assigned to the same class. To avoid this, MuSLA (Kingetsu et al, 2022) employs mutual information of the sample $X$ and the corresponding prediction $\hat{Y}$,
@@
In entropy minimization, the gradient can often be dominated by low-confidence predictions. However, the cross-entropy loss can be too strict when it comes to predicted labels, leading to incorrect updates of the model even when there is a single wrong prediction. To address this trade-off, a new approach, Soft Likelihood Ratio (SLR) loss, is proposed by Mummadi et al (2021) and further employed by Marsden et al (2024). This approach emphasizes predicted classes while considering the issue raised in MuSLA:
$\quad\mathcal{L}_{SLR}(\hat{y}_{ti})=-\sum_{c}w_{ti}\hat{y}_{tic}\log\left(\frac{\hat{y}_{tic}}{\sum_{j\neq c}\hat{y}_{tij}}\right),$ (7)
where $\hat{y}_{ti}$ is the softmax probability for the i-th test sample at time step t. $w_{ti}$ is a ROID-only weight. It is a combination of diversity, considering the similarity between the recent trend of the model’s prediction and the current model output, and certainty, which is the negative entropy of the output. Here, suppose the output confidence for class $c$ is small. In that case, the loss calculation will be reweighted by the summation over all other predictions $\sum_{j\neq c}\hat{y}_{tij}$ in the denominator, thus relaxing the focus on low-confidence classes.
The cooperation between a teacher and a student is another possible solution to optimize prediction confidence with reliability. Here, the teacher is usually the moving averaged model of interest across iterations. CoTTA (Wang et al, 2022a) uses the Softmax prediction from the teacher to supervise the Softmax predictions from the student under the cross-entropy loss.
@@
A cross-entropy loss helps minimize model prediction uncertainty but may fail to give consistent uncertainty scores under different augmentations, etc. This problem is more critical when the teacher-student mechanism is not used. To address this problem, MEMO (Zhang et al, 2022) computes the average prediction across multiple augmentations for each test sample and then minimizes the entropy of the marginal output distribution over augmentations:
@@
To encourage consistency against smaller perturbations, Marsden et al (2024) proposes a consistency loss based on symmetric cross-entropy loss (SCE) (Wang et al, 2019):
$\mathcal{L}_{SCE}(\hat{y}_{ti};y_{ti})=$ (10)
$-\frac{w_{ti}^{\prime}}{2}\left(\sum_{c=1}^{C}\hat{y}_{tic}\log\tilde{y}_{tic}+\sum_{c=1}^{C}\tilde{y}_{tic}\log\hat{y}_{tic}\right).$
This is enabled by promoting similar outputs between test images which have been identified as certain and diverse and an augmented view of them. Here, $\hat{y}_{ti}$ is the softmax probability for the i-th test sample at time step t. $\tilde{y}_{t}$ is the softmax probability of the augmented view. $w_{ti}^{\prime}$ is the weight of the augmented view as in Eq. (7).
@@
This term aims to identify a weight perturbation $\epsilon$ within a Euclidean ball of radius $\rho$ that maximizes entropy. It quantifies sharpness by measuring the maximal change in entropy between $\Theta$ and $\Theta+\rho$. As such, Eq. (12) jointly minimizing entropy and its sharpness. Gong et al (2023) uses same idea for their optimization.
@@ 3.1.5 Other Approaches
To prevent loss of generalization and catastrophic forgetting, weight ensembling integrates a solution. ROID (Marsden et al, 2024) continuously ensemble the weights of the initial source model and the weights of the current model at time step $t$ using moving average, allowing for partial retention of the source knowledge. This approach is akin to the parameter reset strategy in CoTTA, commonly employed in continual adaptation tasks. Additionally, to address temporal correlation and class-imbalanced issues during adaptation, it introduces the idea of prior correction. The intuition is that if the class distribution within a batch tends to be uniform, strong smoothing is applied to ensure that no class is favored. This is indicated by the sample mean over the current softmax prediction $\hat{p}_{t}$. Thus, the smoothing scheme is defined as:
$\quad\bar{p}_{t}=\frac{\hat{p}_{t}+\gamma}{1+\gamma N_{c}},$ (13)
where $N_{c}$ denotes the number of classes and $\gamma$ is an adaptive smoothing factor.
@@ 3.2.1 Data Augmentation
To ensure consistent and reliable predictions, ROID (Marsden et al, 2024) applies augmentation for prediction consistency by SCE. PAD (Wu et al, 2021) employs multiple augmentations of a single test sample for majority voting. This is grounded in the belief that if the majority of the augmented views yield the same prediction, it is likely to be correct, as it demonstrates insensitivity to variations in style. Instead, TTPR (Sivaprasad et al, 2021) adopts KL divergence to achieve consistent predictions. For every test sample, it generates three augmented versions. The model is then refined by aligning the average prediction across these augmented views with the prediction for each view. Another approach is MEMO, which uses AugMix (Hendrycks et al, 2020) for test images. For a test data point, a range (usually $32$ or $64$) of augmentations from the AugMix pool $\mathcal{A}$ is generated to make consistent predictions.
@@ 3.2.2 Memory Bank
Contrastive learning is well-suited for OTTA, as discussed in 3.1.3. However, this approach can be challenging for online learning, especially when it pushes away feature representations of data from the same class. Unlike conventional methods where the feature space can be revisited multiple times, AdaContrast (Chen et al, 2022) offers an innovative solution. It keeps all previously encountered key features and pseudo-labels in a memory queue to avoid forming ‘push-away’ pairs from the same class. This method speeds up the learning process and reduces the risk of error accumulation in data from the same class, thereby improving the efficiency and precision of the learning process.
@@ 3.3.1 Module Addition
In the case of continual adaptation, promptly detecting and adapting to changes in data distribution is inevitable to deal with catastrophic forgetting and the accumulation of errors. To realize this, ViDA (Liu et al, 2023b) utilizes the idea of low/high-rank feature cooperation. Low-rank features retain general knowledge, while high-rank features better capture distribution changes. To obtain these features, the authors introduce two adapter modules correspondingly parallel to the linear layers (if the backbone model is ViT).
Additionally, since distribution changes in continual OTTA are unpredictable, strategically combining low/high-rank information is crucial. Here, the authors use MC dropout (Gal and Ghahramani, 2016) to assess model prediction uncertainty about input $x$. This uncertainty is then used to adjust the weight given to each feature. Intuitively, if the model is uncertain about a sample, the weight of domain-specific knowledge (high-rank feature) is increased, and conversely, the weight of domain-shared knowledge (low-rank feature) is increased. This helps the model dynamically recognize distribution changes while preserving its decision-making capabilities.
@@ 3.3.2 Module Substitution
Module substitution typically refers to swapping an existing module in a model with a new one. The commonly used techniques are about:
@@ 3.3.3 Prompt-based Method
The rise of vision-language models (VLMs) demonstrates their remarkable ability in zero-shot generalization. One typical example is CLIP (Radford et al, 2021), which calculates the cosine similarity between text and image features to enable zero-shot inference. To obtain the text feature, a set of predefined prompts is required to combine with the semantic class candidates. [^1] To ensure a clear understanding of this new direction, we merge all the details of the three methods into this section. However, you can check their components in the timeline tables.
Borrows the idea of prompting from VLMs. “Decorate the Newcomers” (DN) (Gan et al, 2023) employs prompts as supplementary information added onto the image input. To infuse the prompts with relevant information, it employs a student-teacher framework in conjunction with a frozen source pre-trained model to capture both domain-specific and domain-agnostic prompts. For acquiring domain-specific knowledge, it optimizes the cross-entropy loss between the outputs of the teacher and student models. Additionally, DN introduces a parameter insensitivity loss to mitigate the impact of parameters prone to domain shifts. This strategy aims to ensure that the updated parameters, which are less sensitive to domain variations, effectively retain domain-agnostic knowledge. Through this approach, DN balances learning new, domain-specific information while preserving crucial, domain-general knowledge.
VLMs often underperform for domain-specific data. While attempting to address this, traditional fine-tuning strategies typically compromise the model’s generalization power by altering its parameters. In response, with the idea of online adaptation, Test Time Prompt Tuning emerges as a solution. Deviating from the conventional methods, it fine-tunes the prompt, adjusting only the context of the model’s input, thus preserving the generalization power of the model. One representative is TPT (Shu et al, 2022). It generates $N$ randomly augmented views of each test image and updates the prompting parameter by minimizing the entropy of the averaged prediction probability distribution. Additionally, a confidence selection strategy is proposed to filter out the output with high entropy to avoid noisy updating brought by unconfident samples. By updating the learnable parameter of the prompt, it could be easier to adapt the model to the new, unseen domains.
((a))
((b))
((c))
@@ 3.3.4 Summary
4 Empirical Studies
Existing OTTA methods predominantly use WideResNet (Zagoruyko and Komodakis, 2016) or ResNet (He et al, 2016) for experiments, overlooking the evolution of backbones in recent years. In this study, we explore the possibility of decoupling OTTA methods from their conventional CNN backbones. We specifically focus on adapting these methods to the Vision Transformer model (Dosovitskiy et al, 2021). Our work presents strategies for adapting methods, initially developed for CNNs, to work effectively with ViT architectures, thereby examining their flexibility under backbone changes.
Baselines. We evaluate eight OTTA methods, using a standardized testing protocol for fair comparison. We use six diverse datasets: three corrupted datasets (CIFAR-10-C, CIFAR-100-C, and ImageNet-C), two real-world shifted datasets (CIFAR-10.1 and OfficeHome), and one comprehensive dataset (CIFAR-10-Warehouse). CIFAR-10-Warehouse plays a pivotal role in our evaluation, featuring a broad array of subsets including real-world variations from different search engines with different colors and images created through the diffusion model. Specifically, we use the Google split in CIFAR-10-Warehouse to assess the capability of OTTA model to handle color shifts alongside mixed object styles. Furthermore, we evaluate the Diffusion split to assess the effectiveness of OTTA on artificially generated image samples, which have become increasingly popular in recent years.
@@ 4.1 Implementation Details
Optimization details. We use PyTorch for implementation on NVIDIA RTX A6000. The foundational backbone for all approaches is ViT-base-patch16-224 (Dosovitskiy et al, 2021) [^2] https://github.com/huggingface/pytorch-image-models. When using CIFAR-10-C, CIFAR-10.1, and CIFAR-10-Warehouse as target domains, we train the source model on CIFAR-10 with $8,000$ iterations, including a warm-up phase spanning 1,600 iterations. The training uses a batch size of 64 and the stochastic gradient descent (SGD) algorithm with a learning rate of 3e-2. We use an identical configuration to train the source model on CIFAR-100, with an extended training duration of $16,000$ iterations and a warm-up period spanning $4,000$ iterations. The source model on the ImageNet-1k dataset is acquired from the Timm repository 333vit_base_patch16_224.orig_in21k_ft_in1k. For the OfficeHome dataset, we train the source model on the clipart domain for $3,500$ with $200$ iterations of warmup. Additionally, we apply basic data augmentation techniques, including random resizing and cropping, across all methods. We use the common optimization setup, employing the Adam optimizer with a momentum term $\beta$ of $0.9$ and a learning rate of 1e-3. Resizing and cropping techniques are applied as a default preprocessing step for all datasets. Then, an input normalization $(0.5,0.5,0.5)$ is adopted to mitigate potential performance fluctuations arising from external factors beyond the algorithm’s core operations.
@@
Figure 8: Comparison of the OTTA performance on the CIFAR-10-C (at severity level $5$) and CIFAR-10.1 datasets with ViT-base-patch16-224. The top and bottom plots show the experiments conducted with batch size $16$ and $1$ based on the LayerNorm updating strategy, respectively. The dotted line represents the source-only performance.
@@
Component substitution. To adapt core methods for use with the Vision Transformer (ViT), we develop a series of strategies:
  • Switch to LayerNorm: In light of the absence of the BatchNorm layer in ViT, it is natural that all BatchNorm-based strategies in the original implementations would be automatically moved onto LayerNorm.
@@
  • Disregard BatchNorm mixup: Removing statistic mixup strategy originally designed for BatchNorm-based methods because LayerNorm is designed to normalize each data point independently.
@@
  • Sample Embedding Changes: For OTTA methods that rely on feature representations, an effective solution is to use the class embedding (e.g., the first dimension of the ViT model feature) as the image feature.
@@
  • Pruning Incompatible Components: Any elements incompatible with the ViT framework are identified and removed.
@@
Above are useful techniques for integrating important OTTA methods into ViT. Note that these solutions are not only limited to the OTTA methods. Rather, they can be viewed as a broader set of guidelines that can be applied where there is a need for changing backbone architectures.
Baselines: We select eight methods that represent the eight OTTA categories mentioned in this paper, respectively. They include:
@@
    Tent: [optimization-based] A fundamental OTTA method rooted in BatchNorm updates for both statistics and affine parameters with entropy minimization. To reproduce it on ViTs, we replace its BatchNorm updates with a LayerNorm updating strategy.
@@
    CoTTA [optimization &\&& data] uses the mean-teacher model, where the teacher model is updated by the moving average of the student. During adaptation, soft entropy is optimized and combined with selective augmentation. For each iteration, it also applies parameter reset (i.e., partially reset the model parameter to the source pre-trained version). While it requires updating the entire student network (i.e., CoTTA-ALL), we further assess CoTTA-LN that only update LayerNorm on its student model. We also deconstruct its parameter reset strategy, resulting in another two variants: updating LayerNorm without parameter reset (CoTTA-LN), and full network updating without parameter reset (CoTTA-ALL).
@@
    SAR [optimization-based] follows the same strategy as Tent while using sharpness-aware minimization (SAM) Foret et al (2021) for flat minima during optimization.
@@
    Conjugate-PL: [optimization-based] As the source model is optimized by the cross-entropy loss, this method is called Conj-CE. It is similar to Tent but allows the model to interact with the data twice for each iteration: once for updating LayerNorm (just like Tent) and another for prediction.
@@
    MEMO: [optimization &\&& data] For each test sample, MEMO applies various data augmentations and adjusts the model parameters to minimize the entropy across the model marginal output distributions from these augmentations. To ensure consistency and avoid unexpected performance fluctuations, we omit all data normalization processes from its set of augmentations. At the same time, we evaluate the performance of MEMO on two versions, LayerNorm update and full update, resulting in MEMO-LN and MEMO-ALL correspondingly.
@@
    RoTTA [optimization, data &\&& model] uses a memory bank to store class-balanced data, considering uncertainty and the “age” of each saved data sample. It also introduces a time-aware reweighting strategy to enhance adaptation stability. For evaluation, given that BatchNorm is not compatible with ViTs, we exclude the RBN module from its implementation.
@@
    TAST [optimization, data &\&& model] integrates multiple adaptation modules into the source pre-trained model. Based on BatchEnsemble Wen et al (2020), these modules are appended to the top of the pre-trained feature extractor. The adaptation modules are then updated multiple times independently by merging their averaged results with the corresponding pseudo-labels for a batch of data. To accommodate the ViT architecture (especially ViT-base), we use the class embedding from the first dimension as the feature representation.
  8. 8.
    ROID [^4] For fair comparison, we do not use gradient accumulation. [optimization &\&& data] is a method for not only typical OTTA but also universal TTA, capable of handling temporal correlation and domain non-stationarity. It incorporates weighted SLR loss with the Symmetric cross-entropy loss, alongside weight ensemble and prior correction mechanisms to ensure efficacy in complex scenarios.
Despite the broad range of available OTTA methods, we believe that a detailed examination of this carefully selected subset will yield valuable insights. Our empirical study is designed to explore the following key questions.
4.2 Does OTTA still work with ViT?
To assess the transferability and adaptability of the selected OTTA methods, we compare them against the source-only baseline (i.e., direct inference) on vision transformers. For consistency and controlled variable comparison, each bar chart included in our analysis is plotted based on the LayerNorm updating strategy.